OpenStudy (anonymous):

If x[n]=0.99......9(n decimal places), what is the lim as n approaces infinity?

OpenStudy (shubhamsrg):

any guesses?

OpenStudy (anonymous):

1?

OpenStudy (skullpatrol):

Good guess :-)

OpenStudy (anonymous):

wow, that was totally a guess. but i can see why it makes sense

OpenStudy (mathmate):

:)

OpenStudy (anonymous):

but can someone still expain it to me :)

OpenStudy (skullpatrol):

$$0.999...=1$$look at $$\frac{1}{3}+\frac{2}{3} =1$$ $$0.333... + 0.666...= ?$$

OpenStudy (anonymous):

0.999.......

OpenStudy (skullpatrol):

$$\frac{1}{3}=0.333...$$ $$\frac{2}{3}=0.666...$$

OpenStudy (anonymous):

if number of nines approaces infinity \[x=0.999...\]\[10x=9.999...\]\[10x-x=9.999...-0.999...=9\]\[9x=9\]\[x=1\]

OpenStudy (mathmate):

Or let x=0.99999... then 10x=9.99999... 10x-x=9.99999..-0.99999... =9 9x=9 => x=1